OpenStudy (anonymous):

Use DeMoivre's theorem to find the indicated power of (5+5i)^3? (rectangular form)

OpenStudy (anonymous):

first, express the situation: \[ 5+5i=r(\cos\theta+i\sin\theta) \]

OpenStudy (anonymous):

what do we get for "r" and theta?

OpenStudy (anonymous):

r=5 and I am not quite sure how to find theta...is it tan(x)=b/a?